Thirty-seven percent of a number is 28. Find the number to the nearest tenth
A. 97.1

B. 103.6

C. 75.7

D. 132.1


Answer by JimboP1977(311)

let x be the number.
To find 37 per cent of the number we need to divide by 100 and then times by 37.
(x/100)*37 = 28
We need to rearrange to find x. Firstly we need to divide both sides by 37.
(x/100) = 28/37
Then we need to times both sides by 100.
x = 2800/37
x = 75.676 (to 3 decimal places)
Therefore the answer is C.
Does that make any sense? Please get back to me.